OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x – 3y = –12

OpenStudy (anonymous):

x int is found by setting y = 0 and solving for x y int is found by setting x = 0 and solving for y can you try>

OpenStudy (anonymous):

x= (-6, 0) y= (0,4)

OpenStudy (anonymous):

Always always always get these line eqns in the form of y=mx +b. In other words, get y on the left side, and the rest of the stuff on the right side. -3y = -2x -12 y=2x/3 + 12/3 (b=12/3 =4 which is the y-intercept.) The x-intercept occurs when y=0. So, 0 = 2x/3 +4 -2x/3 = 4 x = 4(-3/2) x = -6